OpenStudy (anonymous):

What conic does the equation y^2-x^2=4 represent?

OpenStudy (dls):

Hyperbola.

OpenStudy (anonymous):

@DLS How do you know?

OpenStudy (anonymous):

I thought it would have been a circle

OpenStudy (dls):

\[\LARGE \frac{y^2}{4}-\frac{x^2}{4}=1\]

OpenStudy (dls):

This is the general form of equation for a hyperbola.. For circle it is,(x-a)^2+(y-a)^2=1

OpenStudy (anonymous):

So in order to find what it is, you divide the sides by the 4? @DLS

OpenStudy (dls):

Yes,here.

OpenStudy (anonymous):

Is that how you find out which conics is which? Simply divide what is whatever the number is to both sides?

OpenStudy (anonymous):

@DLS

OpenStudy (dls):

Depends.

OpenStudy (dls):

You have to observe and have an idea about all the conics.